JoelTheBoss (joel_the_boss):

Evaluate, \(log_{2}~\frac{1}{64}\)

JoelTheBoss (joel_the_boss):

Possible answer choices, −6 −32 6 32

JoelTheBoss (joel_the_boss):

@ganeshie8 @mathslover Can you guys help? Or busy, at the moment?

mathslover (mathslover):

Write 64 in terms of 2^n what will be n?

JoelTheBoss (joel_the_boss):

I really don't know this is the first question, were starting to learn about it today, and my teacher doesn't explain anything. :/

mathslover (mathslover):

Okay so, tell me one thing. Can I write 64 as 2^6 ?

OpenStudy (ribhu):

@Joel_the_boss learn a little bit logarithm then, at least simple properties

JoelTheBoss (joel_the_boss):

Yes you can , well I think :/

mathslover (mathslover):

I will write 64 as \(\color{blue}{2 \times 2 \times 2 \times 2 \times 2 \times 2 }\) And if I calculate for the terms in blue, I get, 64. So, I can write 64 as : \(2^6\) Yes?

OpenStudy (ribhu):

so 1/64 can be written as 2^-6 as simple as that @mathslover nicely explained good work

JoelTheBoss (joel_the_boss):

Yes,

mathslover (mathslover):

Okay Joel So, we have : \(\log_2 \cfrac{1}{64} \) as \(\log_2 \cfrac{1}{2^6}\) RIght?

JoelTheBoss (joel_the_boss):

Alright I see how you got that, know what to do? :/

mathslover (mathslover):

Now, I can write \(\cfrac{1}{2^6} \) as \(2^{-6}\) So, I have : \(\log_2 2^{-6} \) Getting it?

JoelTheBoss (joel_the_boss):

ehh, kinda I see what your doing and understand nut this whole area is unknown to me. :(

JoelTheBoss (joel_the_boss):

What*

mathslover (mathslover):

Then I guess, I will have to agree with what @ribhu suggested. Go through with some of the basics of Logarithm and then try to do this question.

mathslover (mathslover):

Here is a link : http://www.purplemath.com/modules/logrules.htm Good Luck! :)

ganeshie8 (ganeshie8):

logarithm is same as exponent

JoelTheBoss (joel_the_boss):

Alright thanks both of you! Yourr some awesome people! @mathslover Thanks a lot. @ribhu Thanks for the help! Nice to meet . :) @ganeshie8 Come in early and steals the show. :P

ganeshie8 (ganeshie8):

\[\large \log_2 \frac{1}{64} = ?\] is same as asking \[\large \frac{1}{64} = 2^{?}\]

ganeshie8 (ganeshie8):

logs are fun, cant resist to answer your first q :)

JoelTheBoss (joel_the_boss):

OH! That's the same thing the @mathslover did but he he did step by step, can it always be done that easily? Any exceptions with other numbers or rules?

mathslover (mathslover):

Haha! Ganeshie8 is a superstar of OpenStudy! Although, listen carefully to what he is explaining as this is pretty good way to learn logarithm. \(\log_a x = y\) then \(x = a^y\)

mathslover (mathslover):

There are some exceptions I guess. You will learn them as soon as you progress with some online study material on Logarithms etc. I'm afraid if I will refer to anything deep here, then you may get confused.

ganeshie8 (ganeshie8):

^

ganeshie8 (ganeshie8):

i think remembering that `logarithm` is exactly same as `exponential function` helps in smoothly learning logarithm properties

mathslover (mathslover):

`correct` :D

JoelTheBoss (joel_the_boss):

No its fine guys, thanks for all the help! :) Lol You guys are the reason I love OS, teach me more than the getting paid to teach. @ganeshie8 Yes I see the similarities. Thanks a lot guys. Ill make sure to tag you three for help xD

ganeshie8 (ganeshie8):

\[logarithm ~= ~exponent\] chant that couple of times ;) if you're a visual learner like me, below is the best video on logs https://www.khanacademy.org/math/algebra2/logarithms-tutorial/logarithm_basics/v/logarithms

JoelTheBoss (joel_the_boss):

ok solved it and got -6, was that right? I simplified then expanded and after that multiplied. Was I right?

ganeshie8 (ganeshie8):

\[\huge\color{Red}{\checkmark}\]

JoelTheBoss (joel_the_boss):

ok thanks :D